OpenStudy (anonymous):

f(x + 1/x) = x^2 + 1/x^2 find f(x)?

OpenStudy (anonymous):

x + 1/x = y x = y -1/x f(y) = (y- 1/x)^2 - 1/(y - 1/x)^2

OpenStudy (anonymous):

Am i correct or is there any other way...??

hartnn (hartnn):

not correct

hartnn (hartnn):

u need to isolate x in x + 1/x = y

hartnn (hartnn):

or

hartnn (hartnn):

x^2+1/x^2 = (x+1/x)^2 - 2 so .... ?

hartnn (hartnn):

f(x + 1/x) = x^2 + 1/x^2 = (x+1/x)^2-2 f(x) = x^2-2

hartnn (hartnn):

got this ?

OpenStudy (anonymous):

Yup.....thxxx...

hartnn (hartnn):

welcome ^_^